Two motorcycles are traveling toward each other on the same road, 200 miles apart. They meet in 2 hours. One is traveling 8mph faster than the other. What is the speed of each motorcycle?
Answer by Alan3354(69443)

Two motorcycles are traveling toward each other on the same road, 200 miles apart. They meet in 2 hours. One is traveling 8mph faster than the other. What is the speed of each motorcycle?
---------------
200/2 = 100 mi/hr
---
x + y = 100
x - y = 8
--------------------- Add